### v1

# metaTextGrad: Automatically optimizing language model optimizers

Abstract
Large language models (LLMs) are increasingly used in learning algorithms, evaluations, and optimization tasks. Recent studies have shown that using LLM-based optimizers to automatically optimize model prompts, demonstrations, predictions themselves, or other components can significantly enhance the performance of AI systems, as demonstrated by frameworks such as DSPy and TextGrad. However, optimizers built on language models themselves are usually designed by humans with manual design choices; optimizers themselves are not optimized. Moreover, these optimizers are general purpose by design, to be useful to a broad audience, and are not tailored for specific tasks. To address these challenges, we propose metaTextGrad, which focuses on designing a meta-optimizer to further enhance existing optimizers and align them to be good optimizers for a given task. Our approach consists of two key components: a meta prompt optimizer and a meta structure optimizer. The combination of these two significantly improves performance across multiple benchmarks, achieving an average absolute performance improvement of up to 6% compared to the best baseline.
1 Introduction
Large language models (LLMs) are increasingly used in learning algorithms, optimization, and evaluation tasks (yuksekgonul2024textgrad, ; zheng2024judging, ; yang2024large, ; khattab2024dspy, ; madaan2023selfrefineiterativerefinementselffeedback, ). However, algorithms that incorporate LLMs often face significant challenges. Firstly, many of these algorithms are still hand-crafted to a considerable extent, requiring substantial human expertise and effort to design and implement effectively. Second, LLMs are notably sensitive to the specific wording and structure of their instructions (chern2024can, ), making it tedious to improve them effectively to be used in learning algorithms.
Figure 1: Illustration of the meta-optimization process. A meta-optimizer optimizes LLM optimizers by aligning them with specific tasks through task interaction while leveraging the strengths of different optimizers to propose a more effective optimizer.
Many studies have explored prompt optimization approaches to automatically design better prompts and enhance the performance of LLMs. For example, algorithms such as OPRO (yang2024large, ), MIPRO (opsahlong2024optimizinginstructionsdemonstrationsmultistage, ), and TextGrad (yuksekgonul2024textgrad, ) have introduced optimizers based on LLMs that support automatic prompt optimization. However, these optimizers are often fixed, applying the same optimization strategies with the same optimizer prompts independent of the task, lacking a process to align with the specific task. Importantly, these optimizers are general purpose by design, to be useful to many different downstream tasks and be used by large user bases. Furthermore, different optimizers likely excel at different types of optimization tasks; thus, achieving the best of all optimizers introduces either a choice to be made among them or a strategy to ensemble them. There is no method to automatically design and improve the optimizers to call based on the characteristics of a task.
In this work, we explore how to automatically optimize both the structure and the prompt of optimizers. Specifically, we assume that all LLM calls are black-box calls, where the internal states of the LLM, as well as any information such as model gradients, are inaccessible, and only the inputs to and outputs of the LLM are observable. For the task to be optimized, we require only a small training dataset containing some input-output pairs and an evaluation metric the user wants to improve upon.
We propose using a meta-optimizer to automatically optimize the optimizer, with the objective of identifying improved optimizers such that the program optimized by it achieves the best performance on the given task. To this end, we introduce two types of meta-optimization strategies: the meta prompt optimizer and the meta structure optimizer. The meta prompt optimizer focuses on refining the prompts of the LLM optimizers to enhance their effectiveness and better align them with specific tasks. Meanwhile, the meta structure optimizer is designed to automatically determine the optimal combination and sequence of different optimizers or modules based on the characteristics of the task.
By combining these two types of meta-optimizers, we propose metaTextGrad, which integrates both components into a unified framework. Specifically, given a set of input optimizers, metaTextGrad first performs prompt optimization for each optimizer independently. Then, it explores the combination and sequencing of these optimizers to construct a more effective composite optimizer. We conducted experiments on multiple benchmarks, and the results demonstrate that our meta-optimized optimizers consistently outperform the existing ones.
Overall, we summarize our contributions as follows. First, we introduce the concept of meta-optimization, highlighting that existing LLM optimizers often require further task alignment and effective combination through a meta-optimizer to maximize their potential. Second, we develop two types of meta-optimizers: the meta prompt optimizer and the meta structure optimizer. Building on these, we propose metaTextGrad, which integrates both types of meta-optimizers into a unified framework. Lastly, experimental results on multiple benchmarks demonstrate that our method significantly outperforms baseline approaches in both performance and generalization.
2 Problem Statement
Here, we will follow the notation from (opsahlong2024optimizinginstructionsdemonstrationsmultistage, ). Consider an LLM program $\Phi$ that may contain multiple LLM calls forming a pipeline, with each call using a different prompt. The pipeline structure of a program and the prompt corresponding to each LLM call can be learnable and optimized by an LLM optimizer.
The task of the LLM optimizer is to find the optimal program $\Phi$ given a training dataset $\mathcal{D}$ (with pairs of inputs $x$ and outputs $y$), and an evaluation metric $\mu$.
$\Phi^{*}=\mathop{\mathrm{arg\,max}}\limits_{\Phi}\,\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mu(\Phi(x),y)$ (1)
Existing methods attempt to approximate solutions to this optimization problem from different perspectives. Optimizers such as MIPRO (opsahlong2024optimizinginstructionsdemonstrationsmultistage, ), OPRO (yang-etal-2022-re3, ), and TGD (yuksekgonul2024attention, ) aim to find better prompts, while optimizers like ADAS (hu2024automated, ) focus on exploring structures. We unify the existing optimizer algorithms into a general framework, as outlined in Algorithm 1. Importantly, these optimizers are the same type as of $M$, are designed by humans, and their structure and prompts are fixed.
Specifically, each optimizer should be capable of performing four operations:
1. Initialize: Given the training dataset and the program to be optimized, prepare the training scheme.
2. Propose: Suggest improvements to the program, which may involve modifications to the pipeline structure, prompts, or other aspects.
3.Update: Update the current optimal program based on the evaluation results of the improved program and determine the next proposal.
4. ExtractOptimizedProgram: Return the best program found so far.
Algorithm 1 Inner Loop: Optimize $\Phi$ with Optimizer $M$
1:  Input: Optimizer $M$, Initial Program $\Phi$, Max Iterations $I$
2:  Input: Training Data $\mathcal{D}$, Validation Data $\mathcal{D}_{val}$, Metric $\mu$
3:  Output: $\Phi^{*}$, i.e., the optimized version of $\Phi$
4:  $M.\text{Initialize}(\mathcal{D},\Phi)$
5:  for $k\leftarrow 1\textbf{ to }I$ do
6:    $\Phi_{k}\leftarrow M.\text{Propose}()$
7:    $\sigma\leftarrow\frac{1}{|\mathcal{D}_{val}|}\sum_{(x,y)\in\mathcal{D}_{val}}\mu\!\bigl{(}\Phi_{k}(x),y\bigr{)}$
8:    $M.\text{Update}(\Phi_{k},\sigma)$
9:  end for
[…]
11:  return $\bigl{(}\Phi^{*},\sigma^{*}\bigr{)}$
The problem we investigate is how to further enhance existing optimizers and align them to be effective optimizers for a given task, instead of relying on human-written optimizers. To this end, we introduce the concept of a meta-optimizer. Let the optimized program obtained by an optimizer $M$ be denoted as $\Phi^{*}=M.\text{optimize}(\mathcal{D},\Phi)$. The optimization objective of the meta-optimizer is to find improved optimizers such that the program optimized by this optimizer performs better on the corresponding task. In particular, the optimization problem is formalized as:
$M^{*}=\mathop{\mathrm{arg\,max}}\limits_{M}\,\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mu(M.\mathrm{optimize}(\mathcal{D},\Phi)(x),y).$ (2)
In optimization, good initializations often contribute to improved performance, in continuous or discrete optimization alike (li2020influence, ; sutskever2013importance, ). Ideally, the meta-optimization framework should leverage the existing, manually designed optimizers and achieve further improvements. Therefore, the input to the meta-optimizer can include one or more existing optimizers as initialization, ultimately producing an optimized optimizer. The detailed process is illustrated in Algorithm 2.
Algorithm 2 Meta-Optimization of Optimizers
1:  Input: Meta-Optimizer $\widehat{M}$
2:  Input: Max Meta-Iterations $J$, Max Inner-Iterations $I$
3:  Input: Initial optimizers $\{M^{(1)},M^{(2)},\dots,$ $M^{(r)}\}$
4:  Input: Training Data $\mathcal{D}$, Validation Data $\mathcal{D}_{val}$
5:  Input: Metric $\mu$, Initial Program $\Phi$
6:  Output: Optimized optimizer $M^{*}$
7:  $\widehat{M}.\text{Initialize}(\mathcal{D},\{M^{(i)}\}_{i=1}^{r})$
8:  for $j\leftarrow 1\textbf{ to }J$ do
9:    $M_{j}\leftarrow\widehat{M}.\text{Propose}()$
10:    $(\Phi^{*}_{j},\sigma_{j})\leftarrow\text{InnerLoop}\bigl{(}M_{j},\Phi,I,\mathcal{D},\mathcal{D}_{val},\mu)$
11:    $\widehat{M}.\text{Update}\bigl{(}M_{j},\sigma_{j}\bigr{)}$
12:  end for
[…]
14:  return $M^{*}$
Specifically, the meta-optimizer combines different input optimizers to propose new optimizers. The newly proposed optimizers execute the process described in Algorithm 1 to obtain the inner loop optimization results. Based on the performance of the newly designed optimizers, the meta-optimizer updates the current best optimizer and determines the proposal for the next iteration (outer loop).
However, similar to the problem of optimizing a program, finding an optimal optimizer using a meta-optimizer is generally an intractable optimization problem. Therefore, it is necessary to introduce appropriate parameterizations and simplifications to the problem. In our work, we primarily focus on two types of parameterizations for the meta-optimizer:
1. automatically optimizing the prompts in the optimizer to align it with the given task (Meta Prompt Optimizer),
2. automatically optimizing the combination of different optimizers based on the characteristics of the task, forming a new composite optimizer (Meta Structure Optimizer).
Collectively, we can meta-optimize using both parameterizations to obtain better optimizers.
3 metaTextGrad: Automatically Optimizing Language Model Optimizers
In this section, we first introduce the theoretical insights behind the meta optimizer. We then present the two core components of metaTextGrad: the meta prompt optimizer and the meta structure optimizer. Finally, we integrate both components to form the complete metaTextGrad pipeline.
3.1 Theoretical Insight
We present the theoretical motivation for meta optimization, highlighting the importance of aligning the optimizer with the target task. In particular, it is shown that when both the training and test sets are sampled from the same underlying data distribution, an optimizer properly aligned via meta-learning on the training set will, with high probability, produce programs on the test set whose accuracy closely approaches that of the optimal optimizer. In contrast, an optimizer that has not been optimized on the training set lacks such theoretical guarantee.
We begin by introducing the necessary notation. Let $\mu:\mathcal{X}\to[0,1]$ denote the accuracy metric, where $\mathcal{X}$ is the space of natural language outputs. Let $D$ be a distribution over natural language inputs $x$. Let $\Phi_{\theta}$ be an optimizer parameterized by $\theta\in\Theta$. Given input $x$ and access to $\mu$, $\Phi_{\theta}$ makes $T$ zeroth-order queries and returns an optimized output $x_{T}$.
Define the loss of the optimizer on input $x$ as $L(\theta,x)=1-\mu(x_{T})=1-\mu(\Phi_{\theta}(x,T)).$ The population loss is given by $R(\theta)=\mathbb{E}_{x\sim D}[L(\theta,x)].$ Let $S=\{x_{1},\dots,x_{n}\}\sim D^{n}$ denote a dataset sampled from $D$, and define the empirical loss on $S$ as $R_{S}(\theta)=\frac{1}{n}\sum_{i=1}^{n}L(\theta,x_{i}).$
Let $\widehat{\theta}_{S}=\arg\min_{\theta\in\Theta}R_{S}(\theta)$ be the optimizer parameters obtained via meta-learning on the training set $S$, and let $\theta^{\ast}=\arg\min_{\theta\in\Theta}R(\theta)$ be the globally optimal optimizer under distribution $D$. We can now state the following theorem:
Theorem 1.
Let $S_{1}$ and $S_{2}$ be two datasets sampled independently from distribution $D$, with sizes $n$ and $m$, respectively. Then, with probability at least $1-\delta$, the optimizer $\widehat{\theta}_{S_{1}}$ trained on $S_{1}$ satisfies:
$R_{S_{2}}(\widehat{\theta})\leq R(\theta^{\ast})+\sqrt{\tfrac{2\log(6/\delta)}{n}}+\sqrt{\tfrac{\log(6/\delta)}{2m}}.$ (3)
The proof of Theorem 1 is based on Hoeffding’s inequality, and the full derivation is provided in Appendix B. The theorem implies that aligning the optimizer with the task through meta learning can be highly effective.
In contrast, an optimizer $\theta_{0}$ that hasn’t been optimized on the training dataset has no theoretical guarantee. We can still guarantee that $R_{S_{2}}(\theta_{0})$ is similar to $R(\theta_{0})$, but the guarantee says nothing about how large $R(\theta_{0})$ is. If the optimizer is bad on average for this task, it will still be bad on $S_{2}$ with high probability. This provides a theoretical foundation for the design of metaTextGrad.
Figure 2: Illustration of metaTextGrad. metaTextGrad combines a meta prompt optimizer and a meta structure optimizer. Given a set of optimizers, metaTextGrad performs optimization in two steps. First, it individually refines each optimizer by optimizing its prompts to better align with the task. Then, it combines the different prompt-optimized optimizers to construct the final optimizer.
3.2 Meta Prompt Optimizer
LLM optimizers are usually designed by humans with manual design choices. As a result, the optimizers themselves are not optimized or aligned with a given task. We aim to optimize the prompts of LLM optimizers to further enhance their effectiveness and align them with tasks.
3.2.1 Motivation
We begin with an example using the Textual Gradient Descent (TGD) optimizer to illustrate why aligning the optimizer with the task during the optimization process can further improve the effectiveness and efficiency of the optimizer. The purpose of TGD optimizer is to “creatively and critically improve prompts, solutions to problems, code, or any other text-based variable”. While this prompt is general, it only provides a very coarse description of the optimizer’s role as improving variables, leaving key aspects (such as the type of problem to be addressed, the high-level idea of optimization) entirely up to the LLM optimizer to infer and handle within a single query. To be clear, this is by design. Optimizers such as TGD or MIPRO are designed to be general purpose optimizers to serve larger user bases. However, as a result, when addressing specific problems, the prompt of the optimizer is not aligned with the task characteristics, which limits the performance of the optimizer.
Therefore, we aim to optimize this prompt so that the optimizer can address specific types of problems more effectively. This naturally brings to mind the concept of meta-learning, or “learning to learn” (finn2017model, ; schmidhuber1987evolutionary, ). The goal of meta-learning is to train a model to rapidly adapt to new tasks. Similarly, the objective of meta prompt optimizer is to align optimizers with a specific class of tasks.
3.2.2 Implementation Method
Below is a brief introduction to the implementation method of the meta prompt optimizer. The detailed pseudocode can be found in Appendix A.1. To Initialize, the meta prompt optimizer runs a round of optimization on the validation dataset to evaluate and record the initial performance of the optimizer. To Propose, the meta prompt optimizer randomly samples data examples from the training dataset and analyzes the general characteristics of the task type. Based on the current best optimizer prompt, it proposes an improved prompt that is more aligned with the task. To Update, the optimized optimizer undergoes an inner loop optimization test on the validation dataset. If the test results outperform those of the previously best optimizer, the optimizer is updated. To ExtractOptimizedOptimizer, the meta prompt optimizer returns the best optimizer learned so far. Among these steps, Propose is the core of the meta prompt optimizer and the only stage where LLM calls are invoked. For detailed prompts, refer to Appendix D.1.
3.3 Meta Structure Optimizer
A variety of LLM optimizers have been proposed to optimize program structures, prompts, and other components. The meta structure optimizer is designed to automatically optimize the combination and ordering of different optimizers based on the characteristics of the task.
3.3.1 Motivation
Since the introduction of OPRO, many LLM optimizers have been proposed. When designing a meta-optimization approach to create better optimizers, it is useful to leverage the strengths of different optimizers and complementing one another. For instance, some tasks may benefit from first identifying a better LLM program architecture and then exploring how to combine prompts and demonstrations within that structure. Other tasks may require iteratively improving the content of prompts and demonstrations to enhance their quality before exploring their combinations. Based on this idea, we propose the meta structure optimizer to alleviate the need for manually combining different optimizers. Instead, it automatically identifies combinations of optimizers that align with task requirements, achieving better results than using a single optimizer alone.
3.3.2 Implementation Method
Below, we briefly introduce the working principles of the meta structure optimizer. The detailed pseudocode implementation can be found in Appendix A.2.
To Initialize, the meta structure optimizer runs one round of optimization for each input optimizer on the validation dataset, selects the highest score and the corresponding optimizer as the initial best value. To Propose, we provide the meta structure optimizer with a set of reference optimizers. If previously optimized, better-performing optimizers exist, they are also included. Based on this, the meta structure optimizer integrates and proposes an improved optimizer. To Update, the meta structure optimizer evaluates whether the proposed optimizer shows improved performance on the validation dataset and updates the current best optimizer and score. To ExtractOptimizedOptimizer, the meta structure optimizer returns the best optimizer learned so far. The Propose stage is the only stage with LLM calls. For detailed prompts, please refer to Appendix D.2.
3.4 metaTextGrad
In the previous two sections, we introduced the meta prompt optimizer and the meta structure optimizer. metaTextGrad is composed of these two components. Specifically, after receiving a set of input optimizers, metaTextGrad first performs prompt optimization for each optimizer individually and then explores the combination of different optimizers to form a better composite optimizer.
4 Experiment
4.1 Experimental Setup
In this section, we use the existing optimizers from DSPy and TextGrad as baseline methods. We evaluate our approach and the baselines on multiple benchmarks, including BBH (suzgun-etal-2023-challenging, ; srivastava2023imitationgamequantifyingextrapolating, ), MMLU (hendrycks2021measuringmassivemultitasklanguage, ), and GPQA (rein2023gpqagraduatelevelgoogleproofqa, ). To ensure reproducibility, we provide the prompts and structures of the learned programs in Appendix E.
Baselines. We used zero-shot CoT, few-shot CoT (wei2023chainofthoughtpromptingelicitsreasoning, ), self-consistency (selfconsistency, ), best-of-N (bon, ; webb2003physical, ), MIPROv2  (opsahlong2024optimizinginstructionsdemonstrationsmultistage, ), TextGrad TGD (yuksekgonul2024textgrad, ) and ADAS-TG (hu2024automated, ) as baselines. Zero-shot CoT relies on direct chain-of-thought reasoning, whereas MIPRO and TextGrad’s TGD optimizer refine the program’s prompt. ADAS is an algorithm for automatically searching and optimizing the program’s structure. We implemented this algorithm within TextGrad as a baseline, referred to as ADAS-TG.
Benchmarks. We evaluate our method on four widely used, challenging, and diverse benchmarks: BBH Word Sorting, BBH Dyck Languages (suzgun-etal-2023-challenging, ; srivastava2023imitationgamequantifyingextrapolating, ), MMLU Abstract Algebra (hendrycks2021measuringmassivemultitasklanguage, ), and GPQA Diamond (rein2023gpqagraduatelevelgoogleproofqa, ). For detailed settings, please refer to Appendix C. Due to the non-determinism of LLM APIs (non_determinism_gpt4_2023, ), the test accuracy for each benchmark is averaged over five random seeds.
In our experiments, we consider three different levels of LLM calls: (1) LLM calls within the program itself, (2) LLM calls made by the optimizer while refining the program, and (3) LLM calls made by the meta-optimizer when optimizing the optimizer. As shown in Section 4.3, the frequency of these calls decreases significantly across these levels. This hierarchical structure allows for cost-effective resource allocation: the program should use a relatively economical model, the optimizer can leverage a more capable model, and the meta-optimizer should employ the best available model. Consequently, in our experiments, we use GPT-4o-mini for LLM calls within the program, GPT-4o for the MIPROv2 and TGD optimizers, and the o1 model for the structure optimizer and the meta-optimizers.
4.2 Main Results
Method Word Sorting Dyck Languages GPQA Diamond Abstract Algebra Average
Val Test Val Test Val Test Val Test Val Test
Vanilla prompting methods
Zero-shot CoT 0.46 0.55 0.06 0.05 0.32 0.34 0.74 0.70 0.40 0.41
8-shot CoT 0.50 0.52 0.14 0.19 0.32 0.35 0.65 0.71 0.40 0.44
Self-consistency (8) 0.47 0.52 0.10 0.12 0.40 0.42 0.76 0.70 0.43 0.44
Best of N (8) 0.48 0.52 0.14 0.17 0.37 0.40 0.77 0.74 0.44 0.46
TextGrad optimizers
TGD Optimizer 0.54 0.55 0.10 0.10 0.34 0.35 0.76 0.71 0.44 0.43
ADAS-TG 0.58 0.58 0.21 0.16 0.36 0.37 0.75 0.70 0.48 0.45
DSPy optimizers
Zero-shot MIPROv2 0.57 0.55 0.19 0.16 0.43 0.38 0.76 0.77 0.49 0.47
8-shot MIPROv2 0.52 0.57 0.33 0.26 0.37 0.34 0.74 0.65 0.49 0.46
Meta-optimized optimizers
metaTextGrad 0.60 0.65 0.42 0.37 0.45 0.40 0.78 0.71 0.56 0.53
Table 1: Accuracy (%) of GPT-4o-mini on benchmarks. Bold indicates the best result, and underlined text represents the second-best.
As shown in Table 1, we achieve up to an 11% absolute performance improvement across these datasets, with an average performance significantly surpassing existing optimizers. Our method outperforms both TGD and ADAS-TG, which serve as the base candidates for metaTextGrad, across all benchmarks and achieves the best performance on most of them. It’s worth noting that the programs optimized by the meta-optimized optimizers also exhibit interesting properties.
(1) The optimized optimizer is more aligned with specific tasks. For example, in the BBH Dyck Languages task, the generated program includes components such as n type analyzer, and a stack validator, which closely match the nature of the task. In contrast, an unaligned optimizer tends to propose more generic and broadly applicable structures.
(2) The optimized optimizer is more effective in handling finer details. For instance, in multi-step LLM calls, passing both the overall problem and subproblems to each LLM call helps maintain a global understanding throughout the process. This behavior is more frequently observed in the optimized optimizer, whereas the initial optimizer tends to pass only the subproblems to each subpart.
(3) The meta-optimized optimizer generally improves efficiency. Although we allocate six optimization steps per training epoch, we observe that meta-optimized optimizers, due to their stronger task alignment, often achieve significant improvements within the first 1-2 steps. In contrast, other optimizers show more gradual improvements.
Please refer to Appendix E for the optimized programs.
4.3 Cost analysis
In this section, we examine the trade-off between effectiveness and computational cost.
First, we analyze the token usage at different levels within a single epoch of meta optimization. This analysis supports our design choice of using models with different capabilities at different levels. As shown in Table 3, the token usage per optimization epoch on the MMLU Abstract Algebra dataset reveals that higher-level components require significantly fewer tokens than lower-level ones. This justifies our hierarchical design. For comparison, a single round of zero-shot CoT requires approximately 140k tokens, indicating that the overall token consumption of our optimization pipeline remains within a reasonable and practical range.
Level Tokens
Program level $\sim$ 400k
Optimizer level $\sim$ 100k
Meta-optimizer level $\sim$ 2.5k
Table 2: Token analysis on Abstract Algebra.
Model Performance Cost
0-shot CoT (4o-mini) 0.05 0.14\$
Ours (4o-mini) 0.37 0.44\$
0-shot CoT (4o) 0.18 0.52\$
Table 3: Cost analysis on Dyck Languages.
In addition, we evaluate the cost and performance of the zero-shot CoT approach using both GPT-4o-mini and GPT-4o on the BBH Dyck Languages dataset, and compare them to our optimized approach applied to GPT-4o-mini. As shown in Table 3, our method achieves the best performance on BBH Dyck Languages while incurring a lower cost than GPT-4o. This demonstrates that with effective prompt and structure optimization, a smaller model can outperform the zero-shot performance of a larger model. These findings highlight the practical applicability and scalability of our approach.
4.4 Transferability of the optimized optimizer across models and datasets
In this section, we evaluate the transferability of our optimized optimizer across different language models and datasets. As shown in Table 5, our method trained on GPT-4o-mini achieves superior performance compared to unoptimized baselines when evaluated on Claude 3 Haiku. Furthermore, as illustrated in Table 5, our optimizer trained on the GPQA diamond dataset also transfers effectively to the Abstract Algebra dataset. These results demonstrate that our meta-optimized optimizer exhibits strong transferability across models and datasets.
Method (Claude 3 Haiku) Dyck Languages
Val Test
Zero-shot CoT 0.07 0.10
TextGrad optimizers
TGD Optimizer 0.10 0.04
ADAS-TG 0.35 0.34
Optimizers optimized on GPT-4o-mini
metaTextGrad 0.32 0.35
Table 4: Transferability of the optimized optimizer across models.
Method Abstract Algebra
Val Test
Zero-shot CoT 0.74 0.70
TextGrad optimizers
TGD Optimizer 0.76 0.71
ADAS-TG 0.75 0.70
Optimizers optimized on GPQA diamond
metaTextGrad 0.78 0.77
Table 5: Transferability of the optimized optimizer across datasets.
4.5 Analysis of the effectiveness of each meta optimizer
In this section, we analyze the contributions of different components of the proposed meta optimizer. As shown in Table 6, we find that all meta optimizers improve performance on the BBH Dyck Languages benchmark. Optimizers optimized using either method outperform the original optimizer. Among them, the meta prompt optimizer achieves the best improvement when applied to ADAS-TG.
Split 0-shot CoT TGD ADAS-TG TGD (O) ADAS-TG (O) Struct (O) metaTextGrad
Val 0.06 0.10 0.21 0.21 0.42 0.24 0.42
Test 0.05 0.10 0.16 0.24 0.37 0.16 0.37
Table 6: Analysis of the effectiveness of each meta optimizer on Dyck Languages. TGD (O), ADAS-TG (O), and Struct (O) respectively denote the TGD and ADAS-TG optimizers enhanced by the meta prompt optimizer, and the optimizers enhanced by the meta structure optimizer.
Here, metaTextGrad produces the same results as the optimized ADAS-TG because the meta optimizer did not find a better option during the meta structure optimization. So, the optimized ADAS-TG was selected as the best result. Notably, the best meta optimizer varies across benchmarks due to task-specific differences. Despite this, all meta optimizers can effectively enhance the performance of a given optimizer.
5 Related Work
5.1 Prompt optimization
Prompt optimization has proven crucial for improving LLM performance. Initial strategies, such as few-shot learning and in-context learning, demonstrated careful prompt design could significantly boost LLM performance (nori2023can, ). Techniques like chain-of-thought reasoning  (wei2023chainofthoughtpromptingelicitsreasoning, ) and ensemble methods  (lu2024mergeensemblecooperatesurvey, ) also emerged as popular ways to structure prompts for effective problem-solving. However, hand-crafted approaches are limited in their utility. Efforts to automate prompt optimization have led to the development of gradient-based approaches (shin-etal-2020-autoprompt, ; chen2022knowprompt, ). These methods, however, require access to model parameters, which limits their application to open-source models.
To address these constraints, alternative approaches have been proposed that leverage LLMs themselves as prompt optimizers. APE (zhou2022large, ) was among the first to introduce the concept of automatic instruction generation and selection. (ye2023prompt, ) introduced the concept of meta prompt, demonstrating systematically designing meta-prompts can improve prompt quality. DSPy (khattab2024dspy, ) proposed a systematic approach for optimizing LLM programs, integrating structured optimization techniques.
Another line of research explores optimization methods based on textual feedback. ProTeGi (pryzant-etal-2023-automatic, ) first introduced the concept of textual gradients, highlighting that LLMs themselves can serve as a form of a loss function to guide the improvement of LLM programs. Building on this idea, TextGrad (yuksekgonul2024textgrad, ) developed a structured textual gradient descent framework, demonstrating its applicability and effectiveness across multiple disciplines and domains. Concurrently, OPTO opto proposed a related approach in which the optimizer receives an execution trace alongside feedback on the generated output. Semantic gradient descent (seback, ) refines textual gradient descent by enhancing feedback signals.
However, the LLM optimizers proposed in these methods remain fixed during the optimization process, limiting their effectiveness and adaptability. Our approach addresses this limitation by leveraging a meta-optimizer to align LLM optimizers with the task.
5.2 Classical meta-learning
Gradient-based meta-learning algorithms such as Model-Agnostic Meta-Learning (MAML) (maml, ), First-Order MAML, Reptile (reptile, ), and ANIL (anil, ) cast learning to learn as finding a good initialization that can be fine-tuned with only a handful of gradient steps. MAML jointly optimizes across tasks through a bi-level procedure, explicitly encouraging large improvements after one or two inner-loop updates. Reptile shows that a simpler first-order update suffices, while ANIL’s ablation studies reveal that the bulk of MAML’s gains stem from feature reuse rather than rapid weight adaptation, allowing the inner loop to be removed for all but the task-specific head. Despite their successes, these methods utilize fixed optimizers. In contrast, we focus on optimizing LLM-based optimizers rather than classical ones, and we meta-learn the optimizer so that the optimization strategy is aligned with each new task, rather than merely learning a good initialization.
6 Conclusion
In this paper, we propose metaTextGrad, a meta-optimization framework that enhances existing LLM optimizers by aligning them more effectively with tasks. Our method introduces two key components: the meta prompt optimizer, which refines optimizer prompts for better task adaptation, and the meta structure optimizer, which determines the optimal combination of different optimizers. By integrating these two components, metaTextGrad improves the efficiency of LLM optimizers, leading to better performance across a diverse range of benchmarks.
Looking ahead, there are several promising directions for future research. First, even the meta optimizer we proposed can be optimized. In particular, our meta optimizer is designed by ourselves, instead of learned by data, and there are techniques in the meta learning literature that could be adapted to allow this (finn2017model, ; schmidhuber1987evolutionary, ). Second, future work can explore different ways to parameterize the optimizers. This study primarily focuses on refining optimizer prompts and their composition, but meta-optimizers could also be leveraged to automatically enhance the optimization algorithms employed by existing optimizers. We believe that optimizing LLM optimizers is a crucial step toward further improving the performance and task alignment capabilities of LLM-driven systems and has the potential to provide valuable insights to the research community.
Appendix A Implementation Details
In this section, we provide the implementation of the meta optimizers. The LLMCall component generates responses based on the meta optimizer’s prompt. The detailed prompts will be provided in the next section of the appendix.
A.1 Meta Prompt Optimizer
Algorithm 3 Meta Prompt Optimizer
1:  function $\widehat{M}.\text{Initialize}(\mathcal{D},M)$
2:    store$(\mathcal{D})$
3:    $M^{*}\leftarrow M$
4:    $\sigma^{*}\leftarrow 0$
5:    return
6:  end function
7:  function $\widehat{M}.\text{Propose()}$()
8:    question, answer $\sim\mathcal{D}$
9:    proposedOptimizer $\leftarrow\text{LLMCall}(\text{prompt},M^{*},\text{question},\text{answer})$
10:    return proposedOptimizer
11:  end function
12:  function $\widehat{M}.\text{Update}(M,\sigma)$
13:    if $\sigma>\sigma^{*}$ then
14:       $M^{*}\leftarrow M$
15:       $\sigma^{*}\leftarrow\sigma$
16:    end if
17:    return
18:  end function
[…]
20:    return $M^{*}$
21:  end function
A.2 Meta Structure Optimizer
Algorithm 4 Meta Structure Optimizer
1:  function $\widehat{M}.\text{Initialize}(\mathcal{D},\{M^{(i)}\}_{i=1}^{r})$
2:    store$(\mathcal{D},\{M^{(i)}\}_{i=1}^{r})$
3:    $M^{*}\leftarrow\text{None}$
4:    $\sigma^{*}\leftarrow 0$
5:    return
6:  end function
7:  function $\widehat{M}.\text{Propose()}$()
8:    proposedOptimizer $\leftarrow\text{LLMCall}(\text{prompt},M^{*},\{M^{(i)}\}_{i=1}^{r})$
9:    return proposedOptimizer
10:  end function
11:  function $\widehat{M}.\text{Update}(M,\sigma)$
12:    if $\sigma>\sigma^{*}$ then
13:       $M^{*}\leftarrow M$
14:       $\sigma^{*}\leftarrow\sigma$
15:    end if
16:    return
17:  end function
[…]
19:    return $M^{*}$
20:  end function
A.3 Meta Optimizer
For the meta optimizer, the execution process consists of two steps: (1) Each optimizer is individually optimized using the meta prompt optimizer. (2) The optimized optimizers are further refined using the meta structure optimizer.
Appendix B Proof of Theorem 1
Theorem (Restatement of Theorem 1).
Let $S_{1}$ and $S_{2}$ be two datasets sampled independently from distribution $D$, with sizes $n$ and $m$, respectively. Then, with probability at least $1-\delta$, the optimizer $\widehat{\theta}_{S_{1}}$ trained on $S_{1}$ satisfies:
$R_{S_{2}}(\widehat{\theta})\leq R(\theta^{\ast})+\sqrt{\tfrac{2\log(6/\delta)}{n}}+\sqrt{\tfrac{\log(6/\delta)}{2m}}.$
Proof.
According to Hoeffding’s inequality, for any $10<\delta<$ 1:
$\Pr_{S_{1}\sim D^{n}}\Bigl{[}|R(\theta)-R_{S_{1}}(\theta)|>\varepsilon_{n}(\delta)\Bigr{]}\leq\delta,$ (4)
where:
$\varepsilon_{n}(\delta)=\sqrt{\tfrac{\log(2/\delta)}{2n}}$ (5)
Using inequality 4, we can bound the population risk of the learned optimizer. With probability at least $1-\frac{2\delta}{3}$:
$R(\widehat{\theta})\leq R_{S_{1}}(\widehat{\theta})+\varepsilon_{n}(\frac{\delta}{3})\leq R_{S_{1}}(\theta^{\ast})+\varepsilon_{n}(\frac{\delta}{3})\leq R(\theta^{\ast})+2\,\varepsilon_{n}(\frac{\delta}{3}),$ (6)
where $\theta^{\ast}=\arg\min_{\theta}R(\theta)$.
With a union bound, with probability at least $1-\delta$:
$R_{S_{2}}(\widehat{\theta})\leq R(\widehat{\theta})+\varepsilon_{m}(\frac{\delta}{3})\leq R(\theta^{\ast})+2\,\varepsilon_{n}(\frac{\delta}{3})+\varepsilon_{m}(\frac{\delta}{3})$ (7)
In conclusion, with probability at least $1-\delta$ over the draws of $S_{1}$ and $S_{2}$:
$R_{S_{2}}(\widehat{\theta})\leq R(\theta^{\ast})+\sqrt{\tfrac{2\log(6/\delta)}{n}}+\sqrt{\tfrac{\log(6/\delta)}{2m}}$ (8)
Appendix C Details on Experimental Benchmark Setup
BBH.
The BBH task [15, 16] is a challenging benchmark that requires precise reasoning across various tasks. We select two subsets from the BBH benchmark: BBH Word Sorting and BBH Dyck Languages. The BBH Word Sorting task requires the language model to sort a given set of words in order, while the BBH Dyck Languages task involves providing a string composed of various types of brackets and asking the model to determine the characters needed to complete the bracket pairing. We use the same train/validation/test splits as in TextGrad [1] (i.e., 50, 100, and 100 instances for training, validation, and testing) and follow the TextGrad approach by using GPT-4o to evaluate whether the model’s output is correct, based on the predicted and ground truth answers.
MMLU.
The MMLU benchmark [17] evaluates a model’s ability to answer questions across a wide range of scientific disciplines. We selected the MMLU Abstract Algebra dataset and created training, validation, and test sets consisting of 10, 50, and 40 questions, respectively. We assess model accuracy using exact matching, determining correctness by applying a regular expression to check whether the model’s response contains "Answer: [A-D]" and matches the ground truth. However, we observed that the DSPy baselines struggle to adhere to this output format. To accommodate this, we use GPT-4o to evaluate the correctness of responses generated by DSPy-based methods.
GPQA Diamond.
The GPQA Diamond benchmark [18] evaluates the model’s ability to solve graduate-level, Google-proof questions. The benchmark consists of 198 questions, which we split into training, validation, and test datasets containing 30, 100, and 68 questions. Similar to MMLU, we use exact matching to determine response accuracy by applying a regular expression to check whether the model’s answer follows the format "Answer: [A-D]" and matches the correct answer. For the DSPy baselines, we use GPT-4o to evaluate the correctness of DSPy-generated responses.
Appendix D Prompts of Meta Optimizers
In this section, we provide the prompts used in the propose stage of the meta prompt optimizer and the meta structure optimizer. The propose stage is the only part of the meta optimizer that involves an LLM call; all other functions consist of direct numerical updates or result retrieval, as explained in pseudocode in Appendix A.
D.1 Meta Prompt Optimizer
Below, we present the prompt used for the meta prompt optimizer.
Meta Prompt Optimizer # Task Requirement A TextGrad optimizer optimizes a TextGrad pipeline so that the pipeline can generate better outputs based on inputs. A TextGrad pipeline consists of several agents, each of which has a specific role in the optimization process, which is defined by different prompts. You will be given a general task description of an optimizer and the specific task the pipeline aims to solve. Your task is to propose an improved optimizer task description so that the optimizer can better optimize the pipeline for the given task. # Optimizer Code Here is the source code of the optimizer: (just for reference) {optimizer_source_code} # The task of the optimizer Specifically, the pipeline aims to solve this kind of question: {example_set.get_question_type()}. An example of the task is provided here: Question: {example_question} Answer: {example_answer} The LLM optimizer wants to improve a LLM pipeline to solve such kind of problems. # Current Task Description Here is the current task description of the optimizer, which you can improve: {optimizer_prompt} # Your task You should identify what the optimizer should pay attention to in order to improve the {optimizer_type} of the pipeline for solving the given task. Conduct a detailed analysis of the given example, and respond in the following format: “‘json "improved_task_description": "…", # Your improved task description for the optimizer “‘
D.2 Meta Structure Optimizer
Meta Structure Optimizer # Task Requirement A TextGrad optimizer optimizes a TextGrad pipeline so that the pipeline can generate better outputs based on inputs. The structure and prompts of current optimizers are fixed, and you will be given the source code of some optimizers. Your task is to propose an improved optimizer code to better optimize the pipeline. This can be achieved by integrating the reference codes of the given different optimizers and merging them into a new optimizer.
Meta Structure Optimizer (Continued) # Task Details You will be provided with implementations of several optimizers, each annotated with their respective purposes. Carefully read through their functions and implementation details. Your task is to integrate the different implementation approaches to provide a more optimized solution. In this context, the step() function refers to the process of performing a single optimization step. You may notice that different optimizers are suitable for different purposes and should be applied in a specific order. For example, if the pipeline aims to optimize the structure, this optimization will overwrite the prompts of each previously optimized component, so the structure optimization should be executed first. Your implementation of the step() function should take the order into account, and each call to the step() function should only optimize a single aspect of the pipeline because the gradient context is not reusable. Therefore, you need to implement a logic so that the optimizer can use different optimizing strategies in a sequential manner, with each strategy being applied self.epoch times before moving to the next one. {optimizer_prompt} # Code Reminder You should include the following necessary imports at the beginning of the code: import inspect import copy from typing import Type, List, Union from collections import defaultdict from textgrad.variable import Variable from textgrad import logger from textgrad.engine import EngineLM from textgrad.optimizer.optimizer import Optimizer, get_gradient_and_context_text from textgrad.model import Pipeline from textgrad.autograd import FormattedLLMCall from textgrad.config import validate_engine_or_get_default from textgrad.optimizer.optimizer_prompts import construct_tgd_prompt, OPTIMIZER_SYSTEM_PROMPT, PIPELINE_SYSTEM_PREFIX, PIPELINE_SYSTEM_SUFFIX # Note You are required to build on one of the existing optimizers and improve it by integrating features from the others, instead of starting from scratch. You should only include the ImprovedOptimizer (inherited exactly from Optimizer) class in the code, and no other classes or functions. You should implement all the necessary functions and attributes in the ImprovedOptimizer class to ensure that the code can be executed without errors. You should not modify input signatures of the __init__ and step functions when implementing the ImprovedOptimizer class. Please output the complete improved Python code, and make sure to call the improved class ’ImprovedOptimizer’; remember to also include all necessary attributes of the class so that the code can be executed without errors. Do not include any additional comments or unnecessary text in the output.
Appendix E Optimized Programs
To facilitate the reproducibility of our work, we provide the program generated by the optimizer optimized with metaTextGrad.
E.1 BBH Word Sorting Benchmark

### v2

@@
 Algorithm 1 Inner Loop: Optimize $\Phi$ with Optimizer $M$
-1:  Input: Optimizer $M$, Initial Program $\Phi$, Max Iterations $I$
+1:  Input: Optimizer $M$, Initial Program $\Phi$, Max Iterations II
 2:  Input: Training Data $\mathcal{D}$, Validation Data $\mathcal{D}_{val}$, Metric $\mu$
@@
 9:  end for
-[…]
+10:  $(\Phi^{*},\sigma^{*}\bigr)\leftarrow M.\text{ExtractOptimizedProgram}()$
 11:  return $\bigl(\Phi^{*},\sigma^{*}\bigr)$
@@
 1:  Input: Meta-Optimizer $\widehat{M}$
-2:  Input: Max Meta-Iterations $J$, Max Inner-Iterations $I$
+2:  Input: Max Meta-Iterations $J$, Max Inner-Iterations II
 3:  Input: Initial optimizers $\{M^{(1)},M^{(2)},\dots,$ $M^{(r)}\}$
@@
 12:  end for
-[…]
+13:  $M^{*}\leftarrow\widehat{M}.\text{ExtractOptimizedOptimizer}()$
 14:  return $M^{*}$
@@
 In this section, we first introduce the theoretical insights behind the meta optimizer.
-We then present the two core components of metaTextGrad: the meta prompt optimizer and the meta structure optimizer.
-Finally, we integrate both components to form the complete metaTextGrad pipeline.
+Next, we further analyze and illustrate our motivation through concrete examples.
+Finally, we present the metaTextGrad pipeline, which consists of the meta prompt optimizer and the meta structure optimizer.
 3.1 Theoretical Insight
@@
 $R_{S_{2}}(\widehat{\theta})\leq R(\theta^{\ast})+\sqrt{\tfrac{2\log(6/\delta)}{n}}+\sqrt{\tfrac{\log(6/\delta)}{2m}}.$ (3)
-The proof of Theorem 1 is based on Hoeffding’s inequality, and the full derivation is provided in Appendix B.
-The theorem implies that aligning the optimizer with the task through meta learning can be highly effective.
+The proof of Theorem 1 is based on Hoeffding’s inequality, and the full derivation is provided in Appendix B. The theorem highlights the necessity of performing meta-optimization.
 In contrast, an optimizer $\theta_{0}$ that hasn’t been optimized on the training dataset has no theoretical guarantee.
@@
 Then, it combines the different prompt-optimized optimizers to construct the final optimizer.
-3.2 Meta Prompt Optimizer
+3.2 Motivating Example
+Existing optimizers such as TextGrad and DSPy are manually designed by humans with the goal of performing well across a broad distribution of tasks, and they indeed demonstrate strong average performance.
+For example, the prompt used by the TextGrad TGD optimizer is as follows:
+TextGrad TGD optimizer Prompt You are part of an optimization system that improves text (i.e., variable).
+You will be asked to creatively and critically improve prompts, solutions to problems, code, or any other text-based variable.
+As can be seen, this prompt is indeed highly general, with phrasing such as ‘improve prompts, solutions to problems, code, or any other text-based variable’.
+However, in some cases, such general-purpose prompt fails to effectively optimize model performance.
+While the optimizers in both TextGrad and DSPy can obtain learning signals from task-specific evaluators, these signals tend to be noisy and sparse.
+First, when feedback is provided solely in the form of scalar scores rather than textual guidance, it becomes sparse, making optimization substantially more challenging.
+Second, although optimizers such as TextGrad’s TGD can accept textual gradient feedback, such feedback is often highly noisy, making it difficult for the optimizer to learn effectively.
+For example, the evaluator feedback received by the TGD optimizer may look like the following:
+Feedback Received by the TGD optimizer To improve the prompt for the executer and enhance the objective function, consider the following feedback: 1.
+**Explicit Criteria Definition**: The prompt should explicitly instruct the executer to define the criteria for optical activity at the beginning of the response.
+This can prevent ambiguity and ensure that the executer uses the correct scientific principles.
+For example, the prompt could include a directive to "List the criteria for optical activity before analyzing each compound." 2.
+**Data Verification Directive**: Incorporate a step in the prompt that requires the executer to verify the input data against reliable sources.
+This could be phrased as "Cross-check the properties of each compound with a trusted chemical database before proceeding with the analysis." 3.
+**Structured Logical Reasoning**: Encourage a structured approach to reasoning by breaking down the analysis into distinct steps.
+The prompt could suggest a format like "For each compound, first identify chiral centers, then assess symmetry, and finally determine optical activity".
+Feedback Received by the TGD optimizer (continued) 4.
+**Cross-Referencing Encouragement**: … By incorporating these elements into the prompt, the executer can be guided to produce more accurate and reliable responses, thereby improving alignment with the ground truth answer and enhancing the objective function.
+Here, the evaluator feedback includes suggestions such as ‘List the criteria for optical activity before analyzing each compound’, which are specific to a single problem instance.
+Such feedback is clearly noisy.
+Since a generic optimizer relies solely on feedback to guide its updates, it is likely to incorporate such suggestions into the optimized LLM program prompt, which can be detrimental to the overall performance of the program.
+Yet, in reality, we can choose to let the LLM optimizer adapt to a specific task distribution in order to achieve better performance.
+This is because if the distribution of programs generated by the task-specific optimizer is more aligned with the task requirements, the difficulty of finding the optimal program will be significantly reduced, and the impact of noise in the evaluator’s signal will be greatly mitigated, even when the feedback is noisy.
+We illustrate this using the BBH Dyck Languages task as an example, showing that aligning the LLM optimizer to a specific task distribution can lead to improved performance.
+For instance, consider the following task-specific optimizer prompt:
+A Task-specific Optimizer Prompt You are part of an optimization system specialized in improving prompts for bracket matching and sequence completion tasks.
+Your role is to enhance prompts that help solve Dyck language problems, which involve proper nesting and closure of different types of brackets (, <>, ()).
+When improving prompts, focus on these critical aspects: (1) maintaining accurate bracket pair matching, (2) preserving the LIFO (Last In First Out) order of nested structures, (3) handling multiple bracket types simultaneously, and (4) ensuring complete closure of all open brackets.
+You should critically analyze how the prompts can better guide the model to track open brackets, maintain proper nesting order, and systematically complete sequences.
+Consider incorporating pattern recognition strategies and explicit validation rules in the improved prompts.
+Your improvements should lead to more reliable and accurate bracket sequence completions.
+It can be observed that the task-specific optimizer leads to a shift in the distribution of LLM programs it tends to optimize, making it more likely to generate content that aligns with key task requirements, such as producing programs that satisfy the requirement of preserving the LIFO (Last In First Out) order, etc.
+As a result, even if the evaluator feedback is somewhat noisy or sparse, the LLM program optimized by the optimizer can still perform well, and is more likely to generate critical statements such as: Explicitly push each opening symbol onto the stack and pop it when a corresponding closing symbol is encountered.
+After processing each symbol, describe the current state of the stack, focusing on unmatched opening symbols.
+In contrast, if a generic optimizer is used, it becomes difficult to generate effective prompts under noisy or sparse feedback conditions.
+This demonstrates that adapting to a new task distribution is meaningful.
+To avoid adapting to each new task distribution by hand, we meta-learn how to adapt, which is the core motivation of metaTextGrad.
+3.3 Pipeline
+3.3.1 Meta Prompt Optimizer
 LLM optimizers are usually designed by humans with manual design choices.
@@
 We aim to optimize the prompts of LLM optimizers to further enhance their effectiveness and align them with tasks.
-3.2.1 Motivation
-We begin with an example using the Textual Gradient Descent (TGD) optimizer to illustrate why aligning the optimizer with the task during the optimization process can further improve the effectiveness and efficiency of the optimizer.
-The purpose of TGD optimizer is to “creatively and critically improve prompts, solutions to problems, code, or any other text-based variable”.
-While this prompt is general, it only provides a very coarse description of the optimizer’s role as improving variables, leaving key aspects (such as the type of problem to be addressed, the high-level idea of optimization) entirely up to the LLM optimizer to infer and handle within a single query.
-To be clear, this is by design.
-Optimizers such as TGD or MIPRO are designed to be general purpose optimizers to serve larger user bases.
-However, as a result, when addressing specific problems, the prompt of the optimizer is not aligned with the task characteristics, which limits the performance of the optimizer.
-Therefore, we aim to optimize this prompt so that the optimizer can address specific types of problems more effectively.
-This naturally brings to mind the concept of meta-learning, or “learning to learn” (finn2017model, ; schmidhuber1987evolutionary, ).
-The goal of meta-learning is to train a model to rapidly adapt to new tasks.
-Similarly, the objective of meta prompt optimizer is to align optimizers with a specific class of tasks.
-3.2.2 Implementation Method
 Below is a brief introduction to the implementation method of the meta prompt optimizer.
@@
 For detailed prompts, refer to Appendix D.1.
-3.3 Meta Structure Optimizer
+3.3.2 Meta Structure Optimizer
 A variety of LLM optimizers have been proposed to optimize program structures, prompts, and other components.
 The meta structure optimizer is designed to automatically optimize the combination and ordering of different optimizers based on the characteristics of the task.
-3.3.1 Motivation
-Since the introduction of OPRO, many LLM optimizers have been proposed.
-When designing a meta-optimization approach to create better optimizers, it is useful to leverage the strengths of different optimizers and complementing one another.
-For instance, some tasks may benefit from first identifying a better LLM program architecture and then exploring how to combine prompts and demonstrations within that structure.
-Other tasks may require iteratively improving the content of prompts and demonstrations to enhance their quality before exploring their combinations.
-Based on this idea, we propose the meta structure optimizer to alleviate the need for manually combining different optimizers.
-Instead, it automatically identifies combinations of optimizers that align with task requirements, achieving better results than using a single optimizer alone.
-3.3.2 Implementation Method
-Below, we briefly introduce the working principles of the meta structure optimizer. The detailed pseudocode implementation can be found in Appendix A.2.
+Below, we briefly introduce the working principles of the meta structure optimizer.
+The detailed pseudocode implementation can be found in Appendix A.2.
 To Initialize, the meta structure optimizer runs one round of optimization for each input optimizer on the validation dataset, selects the highest score and the corresponding optimizer as the initial best value.
@@
 For detailed prompts, please refer to Appendix D.2.
-3.4 metaTextGrad
+3.3.3 metaTextGrad
 In the previous two sections, we introduced the meta prompt optimizer and the meta structure optimizer.
@@
 In contrast, other optimizers show more gradual improvements.
-Please refer to Appendix E for the optimized programs.
+Please refer to Appendix E for the optimized programs and Appendix  F for the optimized optimizers.
 4.3 Cost analysis
@@
 1:  function $\widehat{M}.\text{Initialize}(\mathcal{D},M)$
-2:    store$(\mathcal{D})$
+[…]
 3:  $M^{*}\leftarrow M$
@@
 6:  end function
-7:  function $\widehat{M}.\text{Propose()}$()
+7:  function $\widehat{M}.\text{Propose()}$
 8:  question, answer $\sim\mathcal{D}$
@@
 18:  end function
-[…]
+19:  function $\widehat{M}.\text{ExtractOptimizedOptimizer}(\text{})$
 20:  return $M^{*}$
@@
 1:  function $\widehat{M}.\text{Initialize}(\mathcal{D},\{M^{(i)}\}_{i=1}^{r})$
-2:    store$(\mathcal{D},\{M^{(i)}\}_{i=1}^{r})$
+[…]
 3:  $M^{*}\leftarrow\text{None}$
@@
 6:  end function
-7:  function $\widehat{M}.\text{Propose()}$()
+7:  function $\widehat{M}.\text{Propose()}$
 8:  proposedOptimizer $\leftarrow\text{LLMCall}(\text{prompt},M^{*},\{M^{(i)}\}_{i=1}^{r})$
@@
 17:  end function
-[…]
+18:  function $\widehat{M}.\text{ExtractOptimizedOptimizer}(\text{})$
 19:  return $M^{*}$
@@
 Proof.
-According to Hoeffding’s inequality, for any $10<\delta<$ 1:
+According to Hoeffding’s inequality, for any $0<\delta<1$:
 $\Pr_{S_{1}\sim D^{n}}\Bigl[|R(\theta)-R_{S_{1}}(\theta)|>\varepsilon_{n}(\delta)\Bigr]\leq\delta,$ (4)
@@
 E.1 BBH Word Sorting Benchmark
+[…]
+E.2 BBH Dyck Languages Benchmark
+[…]
+[…]
+[…]
+E.3 GPQA Diamond Benchmark
+[…]
+E.4 MMLU Abstract Algebra Benchmark
+[…]
+Appendix F Optimized Optimizers
+F.1 Optimized TGD Optimizer Demo
+The TGD optimizer is generic, so its optimized program offers only general guidance like “ensuring that every opening bracket has a corresponding closing bracket,” without task-specific strategies.
+In contrast, the optimized TGD aligns closely with the task, emphasizing ideas like “preserving the LIFO (Last In First Out) order of nested structures.” Since the optimized TGD optimizer is better aligned with the task, its optimized program also effectively implements the LIFO principle, making it more effective.
+This is why a task-specific prompt optimizer is necessary.
+TGD Optimizer prompt You are part of an optimization system that improves text (i.e., variable).
+You will be asked to creatively and critically improve prompts, solutions to problems, code, or any other text-based variable.
+Optimized program prompt You will answer a reasoning question.
+Think step by step, ensuring logical consistency and accuracy.
+Explicitly define the role and rules for each type of bracket in the sequence.
+Verify each step of your reasoning, ensuring that every opening bracket has a corresponding closing bracket.
+Compare your predicted sequence with a known ground truth or expected pattern, and explain any discrepancies.
+If errors are identified, re-evaluate and correct them.
+Use mental visualization techniques to aid in understanding sequences.
+Provide a clear and concise explanation.
+The last line of your response should be of the following format: ’Answer: \$VALUE’ where VALUE is the answer to the question.
+Meta-optimized TGD Optimizer prompt You are part of an optimization system specialized in improving prompts for bracket matching and sequence completion tasks.
+Your role is to enhance prompts that help solve Dyck language problems, which involve proper nesting and closure of different types of brackets ({}, <>, ()).
+When improving prompts, focus on these critical aspects: (1) maintaining accurate bracket pair matching, (2) preserving the LIFO (Last In First Out) order of nested structures, (3) handling multiple bracket types simultaneously, and (4) ensuring complete closure of all open brackets.
+You should critically analyze how the prompts can better guide the model to track open brackets, maintain proper nesting order, and systematically complete sequences.
+Consider incorporating pattern recognition strategies and explicit validation rules in the improved prompts.
+Your improvements should lead to more reliable and accurate bracket sequence completions.
+Optimized program prompt You will answer a reasoning question.
+Focus on identifying and providing only the missing closing symbols needed to complete the sequence.
+Use a stack-based approach to track opening and closing symbols, ensuring proper nesting and closure.
+Explicitly push each opening symbol onto the stack and pop it when a corresponding closing symbol is encountered.
+After processing each symbol, describe the current state of the stack, focusing on unmatched opening symbols.
+Ensure that the order of closing symbols matches the reverse order of unmatched opening symbols as they appear in the stack.
+Verify the sequence by checking each symbol’s nesting and closure order.
+Highlight critical decision points, such as when to pop from the stack or add closing symbols.
+Use visual aids if necessary to represent the stack state and sequence operations.
+After completing your analysis, perform a verification check by comparing the predicted closing symbols with the expected output.
+If discrepancies are detected, revisit previous steps to self-correct.
+Conclude with a summary of the final state of the stack and confirm that all brackets are matched.
+The last line of your response should be of the following format: ’Answer: \$VALUE’ where VALUE is the answer to the question.
+F.2 Optimized ADAS-TG Optimizer Demo
+The original program contains generic components like a planner, reasoner, and synthesizer, lacking task-specific focus.
+In contrast, the optimized version adds tailored modules such as a type analyzer, stack validator, and nesting analyzer, aligning closely with BBH Dyck Languages.
+This illustrates the necessity of a task-specific prompt optimizer.
+ADAS-TG Optimizer prompt (Some general instructions are omitted here.) Please produce only the code for the pipeline, with a more systematic or hierarchical structure if possible.
+Meta-optimized ADAS-TG Optimizer prompt (Some general instructions are omitted here.) Please produce only the code for the pipeline, with the following structural improvements: 1) Implement a stack-based mechanism for tracking open brackets, 2) Create separate components for sequence parsing, bracket matching, and completion generation, 3) Include validation checks for proper nesting and bracket type matching, 4) Ensure systematic handling of different bracket types ([], {}, (), <>), 5) Maintain a hierarchical structure that clearly separates the parsing logic from the completion generation.
+The pipeline should efficiently handle nested sequences while preserving the LIFO (Last In, First Out) order of brackets.
+Appendix G Performance When Using an Open-source Model as the Optimizer
+We further evaluate the generality of our framework by adopting fully open-source models.
+Specifically, we employ the non-thinking variant of Qwen3-8B as the program model and Qwen3-235B-A22B as both the optimizer and meta-optimizer.
+The experiments are conducted on the BBH Dyck Languages benchmark.
+As shown in Table 7, metaTextGrad continues to deliver strong performance.
+Appendix H Performance When Applying to a Challenging Benchmark
+To further assess the adaptability of our approach, we evaluate metaTextGrad on the ARC-AGI benchmark, a task family known for requiring abstract reasoning and compositional generalization.
+Following common ADAS practice, we sample ARC-AGI instances with grid sizes $\leq 5\times 5$, comprising 20 training, 30 validation, and 30 test examples.
+The evaluation reports one-shot success rates, using Claude 3 Haiku as the program model and Claude 3.5 Sonnet as both optimizer and meta-optimizer.
+As presented in Table 7, the results demonstrate that metaTextGrad remains highly competitive in this more demanding setting.
+Method Dyck Languages (Qwen models) ARC-AGI (Challenging Benchmark)
+Val Test Val Test
+Vanilla prompting methods
+Zero-shot CoT 0.27 0.27 0.27 0.23
+8-shot CoT 0.37 0.40 0.03 0.00
+Self-consistency (8) 0.31 0.32 0.30 0.23
+Best of N (8) 0.39 0.41 0.27 0.20
+TextGrad optimizers
+TGD Optimizer 0.69 0.68 0.33 0.33
+ADAS-TG 0.32 0.34 0.28 0.26
+DSPy optimizers
+Zero-shot MIPROv2 0.59 0.50 0.30 0.23
+8-shot MIPROv2 0.57 0.51 0.33 0.03
+Meta-optimized optimizers
+metaTextGrad 0.82 0.77 0.37 0.40
+Table 7: Validation and test accuracy when using open-source models (Qwen models, Dyck Languages) and evaluating on a challenging benchmark (ARC-AGI).
+Bold entries denote the best-performing method within each benchmark.
+Appendix I Potential Limitations, Societal Consequences, and Broader Impacts
+Limitations. While metaTextGrad demonstrates stable performance across various models and benchmarks, we acknowledge the following limitations:
+(1) Although metaTextGrad improves performance on many benchmarks, it may not be effective in scenarios where the base model lacks sufficient task-relevant knowledge or reasoning capabilities.
+For example, GPT-4o-mini struggles on math competition benchmarks such as AIME 2024, where metaTextGrad alone cannot yield significant performance gains.
+(2) The meta-optimizer relies on strong instruction-following and problem analysis capabilities.
+As a result, it currently requires advanced models such as o1 or Claude-3.5-Sonnet.
+Models like Gemini 1.5 Pro do not yet perform adequately.
+However, given the rapid progress in model development, we expect that more models will become suitable for the framework in the near future.
+Societal Consequences. metaTextGrad can have meaningful societal consequences.
+  • Acceleration of domain-specific AI applications. By making it easier to adapt LLMs to specific tasks, metaTextGrad may accelerate the deployment of more reliable AI solutions in other domains.
+  • Risk of automation bias or over-reliance. As optimizers become more autonomous, users might rely on them without fully understanding their behavior or limitations.
+  • Potential misuse for persuasive or manipulative systems. metaTextGrad could be exploited to generate more persuasive outputs.
+Impact Statement.
+This paper presents work whose goal is to advance the field of Machine Learning.
+The paper is solely centered on the methodology itself.
+How it is applied and for what purpose is entirely up to the users.
+There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We clearly state our claims in the abstract and the introduction section. They accurately reflect the methods in Section 3 and the experiments in Section 4.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We mentioned potential limitations in Appendix I.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We provide the full set of assumptions in Section 3.1 and a complete proof in Appendix B.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We clearly describe our methods in Section 3 and Appendix A. The experimental settings are provided in Section 4 and Appendix C.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We provide open code access.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The experimental settings are provided in Section 4 and Appendix C.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Results shown in the experiment section are repeated for 5 times, using different random seeds.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The API is the only compute resource we used. We report the details of API usage in Section 4.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The paper conforms with the NeurIPS code of ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We discuss the broader impacts in Appendix I. The paper is solely centered on the methodology itself. How it is applied and for what purpose is entirely up to the users.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: They are properly credited, mentioned and respected.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: Yes, code is the only new asset. It is well documented and the documentation is provided alongside the code.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79. Justification: The LLM program, optimizer, and meta-optimizer are the core components of the proposed method and are thoroughly discussed in the paper.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.